# On the PM2.5 – Mortality Relationship: A Bayesian Dynamic Model for Spatio-Temporal Confounding

### v1

@@ Abstract
Spatial confounding, often regarded as a major concern in epidemiological studies, relates to the difficulty of recovering the effect of an exposure on an outcome when these variables are associated with unobserved factors. This issue is particularly challenging in spatio-temporal analyses, where it has been less explored so far. To study the effects of air pollution on mortality in Italy, we argue that a model that simultaneously accounts for spatio-temporal confounding and for the non-linear form of the effect of interest is needed. To this end, we propose a Bayesian dynamic generalized linear model, which allows for a non-linear association and for a decomposition of the exposure effect into two components. This decomposition accommodates associations with the outcome at fine and coarse temporal and spatial scales of variation. These features, when combined, allow reducing the spatio-temporal confounding bias and recovering the true shape of the association, as demonstrated through simulation studies. The results from the real-data application indicate that the exposure effect seems to have different magnitudes in different seasons, with peaks in the summer. We hypothesize that this could be due to possible interactions between the exposure variable with air temperature and unmeasured confounders.
Keywords: Air pollution; Bayesian; Dynamic generalized linear model; Mortality; Non-linearity; Spatio-temporal confounding.
@@ 1 Introduction
In this paper, the ultimate goal is to correctly evaluate the short-term association between particulate matter with a diameter less than $2.5\mu m$ (PM2.5) and all-cause mortality. The data span the period between 2018 and 2022 and are available for each health district of two Italian regions (see Section 2). We argue that the assumption of a constant and linear effect can be unrealistic for several reasons. The physical properties and chemical composition of the PM mixture may vary in time and space, as they are influenced by meteorological conditions, human activities, and the morphology of the study region (Chen et al., 2013, Masselot et al., 2022, Peng et al., 2005). Temperature and sunlight, for example, can improve the photochemical reactions that produce secondary aerosols (Chen et al., 2013, Peng et al., 2005). Exceptional events such as sand storms (Chen et al., 2013) and wildfires (Zhou et al., 2021) cause an increase in the concentration of hazardous particles, so they can modify the association with the outcome. Human activities, such as traffic, industrial production, domestic heating, and air conditioning, change throughout the year, so they may also affect the concentration and the composition of pollutants (Chen et al., 2013, Masselot et al., 2022, Peng et al., 2005, Kelly and Fussell, 2012). The physiological response to air pollution may also vary by season, as extreme temperatures may increase a person’s susceptibility to PM effects (Bergmann et al., 2020, Peng et al., 2005, Roberts, 2004, Li et al., 2017). Finally, during the summer people spend more time outdoors or have windows open, and this increased natural ventilation may enhance people’s exposure to air pollutants and reduce the exposure measurement error (Bergmann et al., 2020, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008).
To account for the potential non-linearity of the exposure–outcome relationship, various strategies are available. An exposure-response function is often estimated (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). Alternatively, it would be appropriate to specify a model that incorporates the inherent spatial or temporal variations of the exposure effect (Gelfand et al., 2003). Here, we focus on the temporal evolution of the impact of PM2.5 on human health. By explicitly modeling the temporal variations in the coefficients, we can analyze whether the strength of the relationship between PM2.5 and health outcomes have changed over different time periods. Furthermore, the temporal dynamic of the exposure effect allows investigating the influence of various factors that have emerged since 2020, such as lockdowns, the use of facial masks, physical distancing and social restrictions. These measures have significantly altered people’s habits and interactions with the environment, potentially affecting the relationship between PM2.5 exposure and health outcomes.
@@
A second major challenge in ecological regression models relates to the exposure effect being influenced by several confounders, that is factors associated with both the exposure and the outcome variables. Although information on some of these confounders is available, such as meteorological variables, it is likely that others are unmeasured or difficult to quantify, such as influenza epidemics or changes in population structure (Peng et al., 2006). If the model does not correctly account for unmeasured confounding, the estimator of the exposure effect is biased. We denote this situation as spatio-temporal confounding. More specifically, we extend the notion of spatial confounding (Paciorek, 2010) to the spatio-temporal set up.
@@
Despite the complexity of this problem, just few authors have recently proposed approaches that explicitly account for spatio-temporal confounding. For instance, Reich et al. (2021) discuss extensions of causal inference methods. Prates et al. (2022) propose the transformed multivariate Gaussian Markov random field model (TMGMRF), an extension of the approach by Prates et al. (2015), where Poisson intensities or Bernoulli rates are modeled directly and without the need of a link function. A copula structure captures the spatio-temporal dependence and guarantees the disentanglement between fixed and random effects, hence their model should be able to control for confounding bias. A spatio-temporal linear model is considered by Adin et al. (2023), where the random effects are orthogonalized with respect to the fixed effects: in other words, this work extends the restricted spatial regression model (RSR; Hodges and Reich, 2010, Reich et al., 2006). In all these approaches, however, the exposure–outcome association is always assumed to be constant and linear.
@@
A comprehensive data-driven approach that can capture the temporal dynamics of the exposure effect while simultaneously controlling for spatio-temporal confounding is still lacking. To tackle both problems, we propose a Bayesian DGLM, which has several advantages. First, it is able to approximate any smooth function of time (Gamerman and Lopes, 2006, Granger, 2008), hence the proposed model should not be affected by misspecification bias. Second, the proposed approach can be seen as a generalization of other methods in the literature. For example, the approaches by Chen et al. (2013) and Peng et al. (2005) cannot account for exceptional events such as the outbreak of the SARS-CoV-2 pandemic. In addition to this issue, the models proposed by Dai et al. (2014) and Klompmaker et al. (2023), among others, implicitly assume the possibility of abrupt changes. However, the latter ones are unlikely in practice, because the exposure effect depends, for example, on meteorological conditions and on human activity patterns that change gradually over time (Peng et al., 2005, Bhaskaran et al., 2013). In contrast, the proposed model is able of overcoming both drawbacks, thanks to the ability of DGLMs to capture gradual changes and to emphasize both seasonal and non-systematic variations (Lütkepohl, 2005, Petris et al., 2009, Bitto and Frühwirth-Schnatter, 2019).
A third major advantage of the proposed approach is that, following Greven et al. (2011) and Janes et al. (2007), the exposure effect is decomposed into two orthogonal components, each describing the (non-linear) association with the health outcome at a different scale of variation. This decomposition should help to distinguish the effect of interest from that of unmeasured confounding factors. Finally, the time-varying spatial-specific intercepts in our specification should allow capturing seasonal and trend components due to the unmeasured confounders.
[…]
2 Motivation
@@
This paper analyzes the association between daily PM2.5 concentrations and all-cause mortality counts among the elderly (i.e. people aged 65 or more) in Piemonte and Lombardia regions (Italy) between 2018 and 2022. In this Section, after describing the dataset, we emphasize the importance of capturing the temporal dynamics of the regression coefficients, an aspect that is not adequately addressed by existing approaches.
2.1 Data Description
@@
The data used in this study are retrieved from publicly available sources and full details are given in Section A of the Supplementary Material. Mortality counts are collected from the Italian National Institute of Statistics (ISTAT, 2023a), and are observed on an irregular lattice comprising ${N}=117$ non-overlapping areal units termed health districts (HDs). The study period comprises ${T}=1826$ consecutive days, hence $Y_{it}$ denotes the number of deaths (by any cause) occurred in district $i=1,\dots,{N}$ during day $t=1,\dots,{T}$. The PM2.5 concentrations are used as a measure of the exposure in district $i$ during day $t$, denoted as $X_{it}$, and are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). We also acquire information on temperature and relative humidity, which are potential confounders, from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018). The remote-sensed measurements are spatially and temporally aggregated to match the resolution of the health data.
@@
Figure 2 shows the average daily SMR (on the logarithmic scale) for Lombardia and Piemonte. The seasonal pattern is similar in both regions, with the lowest values in summer, except for the year 2020. Two peaks can be noticed, where the SMR more than doubled: the first one occurred in April 2020 and the other one in November 2020. These are due to the first and second “waves” of the SARS-CoV-2 pandemic, respectively. The first wave was more severe in Lombardia, while the second one was more severe in Piemonte.
@@
Figure 2: Time series of the average daily SMR for Piemonte (top) and Lombardia (bottom) regions, on the log scale. The vertical dashed lines separate each calendar year.
@@
Figure 2 also highlights other two aspects. One is that the highest values of the SMR in 2018 were reached at the beginning of January, suggesting that the influenza epidemic reached its peak before the end of 2017. This is different from the year 2019, where the SMR peaks were reached in February. Another difference is that there was an increase in mortality during summer 2022. Possible reasons may be the occurrence of prolonged heat waves (indeed, the average daily temperature time-series has its absolute maximum in July 2022), or the outbreak of an influenza epidemic.
2.2 Evidence from Competing Approaches
It is assumed that the outcome follows a Poisson distribution, such that, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$:
$Y_{it}|\mu_{it}\stackrel{{\scriptstyle ind}}{{\sim}}\mathcal{P}oi(\mu_{it})\,,$ (1)
where $\mu_{it}$ indicates the mean. Then the linear predictor, $\vartheta_{it}=\ln(\mu_{it})$, is modeled as a function of exposure, $X_{it}$, and measured confounders, denoted as ${}^{m}\mathbf{Z}_{it}=({{}^{m}Z}_{1it},\dots,{{}^{m}Z}_{{p_{m}}it})^{\prime}$, comprising basis functions of meteorological variables, and indicator variables for weekdays, national holidays, and lockdown periods (hence ${p_{m}}=19$; see also Section A of the Supplementary Material). Additional covariates and basis functions can be included in the linear predictor to account for unmeasured confounding (see, e.g., Peng et al., 2006, Peng and Dominici, 2008). Several approaches from the literature are then fitted under the Bayesian paradigm: (i) the Null model, which does not account for unmeasured confounding; (ii) the GLMadj model, which adjusts for unmeasured confounding through a natural cubic spline of time (see, e.g., Peng and Dominici, 2008, Zhou et al., 2021, Masselot et al., 2022); (iii) the JDZ model by Janes et al. (2007), which is similar to the previous approach but further decomposes the exposure effect in two orthogonal parts; (iv) the Dummy model (see, e.g, Dai et al., 2014, Klompmaker et al., 2023), which allows for season-specific exposure effects; and (v) the Periodic model by Peng et al. (2005) and Chen et al. (2013), which assumes that the exposure effect is a linear combination of trigonometric functions. A hierarchical structure is assumed for each of them, with area-specific regression coefficients having the same mean. Further details are provided in Section C of the Supplementary Material.
Figure 3 shows, for each approach, the (overall) estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in PM2.5 levels, and it can be noticed that the estimates differ considerably. The Null model yields the highest percent change on average ($3.92\%$), whereas the JDZ model estimates the lowest percent change on average ($0.47\%$). The models (i)–(iii) assume that the association of interest is constant over time, however this assumption is unlikely for the arguments outlined in Section 1.
Figure 3: Estimated percent change of risk associated with a $10\ \mu g/m^{3}$ increase in the PM2.5 concentrations under several approaches. Null: the model that does not account for confounding; GLMadj: the model with a smooth function of time; Dummy: the model similar to GLMadj, but it also captures season-specific exposure effects; Periodic: the model similar to GLMadj where the exposure effect is assumed to be a periodic function; JDZ: the model by Janes et al. (2007). All models assume area-specific intercepts.
The only two approaches able to capture varying strengths of the association are the Dummy and the Periodic models. They seem to agree on the temporal evolution of the effect, showing peaks in summer and troughs in winter. Further discussion on this seemingly unexpected finding is delayed to Section 6. The Dummy and the Periodic models, however, have some limitations, since both assume that the effect changes within the year, but is the same across years. This is not consistent with the evidence from the data, as the pandemic outbreak may have altered the strength of the association. In fact, consider Figure 4, showing the average daily PM2.5 concentrations for Lombardia and Piemonte regions. Their temporal evolutions correspond only in part to those of the SMR. In particular, the two peaks in the SMR in 2020 are not reflected in the PM2.5 time series. Additionally, it can be noticed from Table B.3 that the overall 2020 annual mean of PM2.5 levels was the lowest among the five years considered in our analysis. In other words, the steep increase in mortality during that year was not caused by air pollution. Instead, it is expected that people were less exposed to fine particulate matter because of the lockdown measures implemented in Italy to contain the spread of the SARS-CoV-2 virus.
Figure 4: Time series of the average daily PM2.5 concentrations for Piemonte and Lombardia regions. The vertical dashed lines separate each calendar year.
@@ 3 The Proposed Model
For $i=1,\dots,{N}$ and $t=1,\dots,{T}$, let $Z_{it}$ denote an unmeasured confounder. We assume that ${{}^{m}\mathbf{Z}}_{it}$ and $Z_{it}$ are spatially and temporally smoother than the exposure, and that the outcome follows the distribution in Equation (1). The linear predictor is then specified as:
$\vartheta_{it}$ $=o_{it}+\beta_{0i}+f(X_{it};\bm{\gamma}_{f})+{{}^{m}\mathbf{Z}}_{it}^{\prime}\bm{\alpha}_{m}+Z_{it}+u_{it}\,,$ (2)
where $o_{it}$ is the offset, $\beta_{0i}$ is an area-specific intercept, $f(\cdot;\cdot)$ is an unknown smooth function parameterized by the vector $\bm{\gamma}_{f}$, $\bm{\alpha}_{m}$ is a ${p_{m}}$-dimensional vector of coefficients, and $u_{it}$ is a spatially and temporally uncorrelated Gaussian random variable, assumed to have zero mean and constant variance, i.e. $u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})$.
@@
$f(X_{it};\bm{\gamma}_{f})=f(\widehat{X_{t}};\bm{\gamma}_{f})+f^{\prime}(\widehat{X_{t}};\bm{\gamma}_{f})(X_{it}-\widehat{X_{t}})+\varpi_{it}\,,$ (3)
where $\widehat{X_{t}}$ denotes the large-scale (or overall) temporal trend, $f^{\prime}(\widehat{X_{t}};\bm{\gamma}_{f})$ is the first derivative of $f(\cdot;\bm{\gamma}_{f})$ with respect to its first argument, evaluated at $\widehat{X_{t}}$, and $\varpi_{it}$ denotes the Taylor’s remainder. The overall temporal trend can be calculated as a spatial average at each $t=1,\dots,{T}$. Alternatively, $X_{it}$ can be regressed on a smooth function of time (see also Janes et al., 2007):
$X_{it}=h(t;df)+\nu_{it}\,,$ (4)
where $df$ are the degrees of freedom of the smooth term, and $\nu_{it}$ is an uncorrelated Gaussian random variable with zero mean and constant variance. The overall temporal trend is then defined as the fitted values obtained from the model (4), i.e. $\widehat{X_{t}}=\widehat{h}(t;df)$.
Equation (3) provides a decomposition of the exposure effect into two components describing large and small (or local) scales of variation. These components are functions of $\widehat{X_{t}}$ and $X_{it}-\widehat{X_{t}}$, respectively. They are orthogonal because $X_{it}-\widehat{X_{t}}$ is the residual term from the regression in Equation (4). Henceforth, the decomposition above should help to disentangle the effect of interest from that of unmeasured confounding factors. To clarify this, consider that $\widehat{X_{t}}$ captures trend and seasonal components of the exposure, therefore it is expected to be highly correlated with $Z_{it}$, given that the latter is smooth, while the correlation between $X_{it}-\widehat{X_{t}}$ and $Z_{it}$ should be close to zero. Therefore, the unmeasured confounder is more likely to affect the association between outcome and exposure at large scales (Greven et al., 2011, Janes et al., 2007), so the first term in the decomposition of $f(X_{it};\bm{\gamma}_{f})$ in Equation (3) might be biased, but it is less likely that the second term, $f^{\prime}(\widehat{X_{t}};\bm{\gamma}_{f})(X_{it}-\widehat{X_{t}})$, is affected by confounding bias.
@@
$\vartheta_{it}=o_{it}+\beta_{0i}+f(\widehat{X_{t}};\bm{\gamma}_{f})+f^{\prime}(\widehat{X_{t}};\bm{\gamma}_{f})(X_{it}-\widehat{X_{t}})+{{}^{m}\mathbf{Z}}_{it}^{\prime}\bm{\alpha}_{m}+Z_{it}+\tilde{u}_{it}\,.$ (5)
@@
The unknown function evaluated at the overall temporal trend of the exposure, $f(\widehat{X_{t}};\bm{\gamma}_{f})$, and its first derivative in Equation (5) can be seen as functions of time, therefore a time-varying coefficients can adequately represent them (see also West and Harrison, 1997, Section 3.1). Therefore, we propose to change the linear predictor in Equation (2) for:
$\vartheta_{it}=o_{it}+\beta_{0it}+\beta_{1t}(X_{it}-\widehat{X_{t}})+{{}^{m}\mathbf{Z}}_{it}^{\prime}\bm{\alpha}_{m}+\tilde{u}_{it}\,,$ (6)
where $\beta_{1t}=f^{\prime}(\widehat{X_{t}};\bm{\gamma}_{f})$ is the time-varying coefficient representing the local association between exposure and outcome (in the sense of Janes et al. (2007)), while $\beta_{0it}$ is a time-varying area-specific intercept,
$\beta_{0it}=\beta_{0i}+f(\widehat{X_{t}};\bm{\gamma}_{f})+Z_{it}\,,$
@@
We now provide two examples of the function $f(\cdot;\cdot)$ and of the resulting time-varying coefficients in Equation (6). These show that our model generalizes other methods in the literature.

### v2

# On the PM2.5 – Mortality Relationship: A Bayesian Model for Spatio-Temporal Confounding

@@ Abstract
In epidemiological studies about the effects of air pollution on public health, it is common practice to incorporate an arbitrary smooth function of time (or space) in time-series (or spatial) data to account for unknown confounding effects. These smooth functions are inherently linked to exposure, thereby increasing the difficulty of assessing the exposure effect (i.e. the concurvity problem). This issue is particularly challenging in spatio-temporal analyses, where it has been less explored so far. To study the effects of fine particulate matter on mortality in Italy, a model that simultaneously accounts for spatio-temporal confounding and for the non-linear form of the effect of interest is needed. Henceforth, we propose a Bayesian spatial dynamic generalized linear model, which allows for non-linear associations and for a decomposition of the exposure effect. This decomposition accommodates associations with the outcome at fine and coarse temporal and spatial scales of variation. These features, when combined, allow reducing the spatio-temporal confounding bias and recovering the shape of the association, as demonstrated through simulation studies. The real-data analysis indicates that the exposure effect shows a clear temporal dynamic, with peaks during summer. We argue that this may be due to interactions of particulate matter concentrations with air temperature and unmeasured confounders.
Keywords: Air pollution, Bayesian, Concurvity, Dynamic generalized linear model, Non-linearity, Spatio-temporal confounding.
@@ 1 Introduction
In this paper, the ultimate goal is to correctly evaluate the short-term association between particulate matter with a diameter less than $2.5\mu m$ (PM2.5) and all-cause mortality. The data span the period between 2018 and 2022 and are available for each health district of two Italian regions (see Section 2). We argue that the commonly adopted assumption of a constant and linear effect, without considering interaction with confounding variables, may be inadequate, given the spatial and temporal variability of PM2.5 and its associated health risks. Specifically: (i) the physical properties and chemical composition of the PM mixture exhibits substantial variation across locations and over time, driven by environmental conditions, human activities, and episodic phenomena such as sandstorms and wildfires (Chen et al., 2013, Kelly and Fussell, 2012, Masselot et al., 2022, Peng et al., 2005, Zhou et al., 2021); (ii) human susceptibility to PM2.5 is not static, but rather modulated by temporally and spatially varying factors, including extreme temperatures, which can exacerbate health risks in specific regions or seasons (Bergmann et al., 2020, Peng et al., 2005, Roberts, 2004, Li et al., 2017); (iii) the relationship between true and measured exposure varies across time and space, with seasonal factors like increased summer ventilation potentially reducing exposure measurement error (Bergmann et al., 2020, Peng et al., 2006, Roberts, 2004, Stafoggia et al., 2008).
To account for the potential non-linearity of the exposure–outcome relationship, various strategies are available. An exposure-response function is often estimated (see, among others, Brunekreef et al., 2021b, Cork et al., 2023, Dominici et al., 2022, Renzi et al., 2021). Alternatively, it would be appropriate to specify a model that incorporates the inherent spatial and temporal variations of the exposure effect (Gelfand et al., 2003). By doing so, we can analyze whether the strength of the relationship between PM2.5 and mortality have changed over different time periods and across health districts. The temporal dynamic of the exposure effect allows investigating the influence of various factors that have emerged since 2020, such as lockdowns, the use of facial masks, physical distancing and social restrictions. Indeed, these measures have significantly altered people’s habits and interactions with the environment. Furthermore, the health districts are characterized, for example, by morphological and climatic differences which may influence the exposure–outcome relationship. Therefore, it is important to allow for spatially varying coefficients in the model.
@@
Spatio-temporally varying coefficients models (STVCMs) are reviewed by Banerjee et al. (2014) and Sahu (2022), among others, and have been applied to various fields, including healthcare (Cai et al., 2013, Song et al., 2020), air pollution and climate studies (Paez et al., 2008, Bakar et al., 2015). This class of models provide a flexible framework to capture smooth space-time variations of the exposure–outcome association.
A second major challenge in ecological regression models relates to the problem of confounding. When data are spatially correlated, this problem has received a lot of attention in recent years — see, e.g., Reich et al. (2006), Paciorek (2010), Thaden and Kneib (2018), Dupont et al. (2022), Marques et al. (2022), Azevedo et al. (2023) and Zaccardi et al. (2025). Gilbert et al. (2023) and Khan and Berrett (2023) argue that multiple definitions of “spatial confounding” co-exist in the literature. In this paper, we extend one of their notions and refer to spatio-temporal confounding as the problematic assessment of the exposure effect due to concurvity, that is the correlation between exposure and an unknown smooth function of space-time that affects the outcome (Gilbert et al., 2023, Peng et al., 2006). In other words, factors that vary on similar spatial and temporal scales as the exposure are considered potential confounders. Although information on some of these confounders is available, such as meteorological variables, it is likely that others are unmeasured or difficult to quantify, such as influenza epidemics or changes in population structure (Peng et al., 2006). Not correctly accounting for unmeasured confounding may lead to an incorrect evaluation of the exposure–outcome relationship (i.e., confounding bias), as the model may not be able to distinguish between the effect of the exposure and that of the confounders.
@@
Despite the complexity of this problem, just few authors have recently proposed approaches that explicitly account for spatio-temporal confounding. For instance, Reich et al. (2021) discuss extensions of causal inference methods. Other works have adopted a “data analysis” perspective (Khan and Berrett, 2023), where confounding results from the collinearity between exposure and structured random effects. Within this strand of research, Prates et al. (2022) propose the transformed multivariate Gaussian Markov random field model (TMGMRF), an extension of the approach by Prates et al. (2015), where Poisson intensities or Bernoulli rates are modeled directly and without the need of a link function. A copula structure captures the spatio-temporal dependence and guarantees the disentanglement between fixed and random effects, hence their model should be able to control for confounding bias. A spatio-temporal linear model is considered by Adin et al. (2023), where the random effects are orthogonalized with respect to the fixed effects: in other words, this work extends the restricted spatial regression model (RSR; Hodges and Reich, 2010, Reich et al., 2006). These approaches, however, are not intended to adjust for unmeasured confounders (Khan and Calder, 2022).
@@
A comprehensive data-driven approach that can capture the temporal dynamics of the exposure effect while simultaneously controlling for spatio-temporal confounding is still lacking. To tackle both problems, we propose, under the Bayesian paradigm, the spatial DGLM for confounding (SDGLMC), which has several advantages. First, spacetime-varying intercept and slope coefficient are introduced in the model specification to account for unmeasured spatio-temporal confounders and non-linear exposure–outcome associations, respectively. Second, since DGLMs are only locally linear, they can approximate any smooth function of time (Gamerman and Lopes, 2006, Granger, 2008), hence the proposed model should not be affected by misspecification bias. Third, SDGLMC can be seen as a generalization of other methods in the literature. For example, the approaches by Chen et al. (2013) and Peng et al. (2005) cannot account for exceptional events such as the outbreak of the SARS-CoV-2 pandemic. The models proposed, among others, by Dai et al. (2014) and Klompmaker et al. (2023) assume that the effect varies between seasons, but is constant within each season. In other terms, the latter works implicitly assume the possibility of abrupt changes in the exposure effect between adjacent seasons, however these are unlikely in practice, because the exposure effect depends, e.g., on meteorological conditions and on human activity patterns that change gradually over time (Peng et al., 2005, Bhaskaran et al., 2013). In contrast, the proposed model can overcome both drawbacks, thanks to the ability of DGLMs to capture gradual changes and to emphasize both seasonal and non-systematic variations (Lütkepohl, 2005, Petris et al., 2009, Bitto and Frühwirth-Schnatter, 2019). Additionally, SDGLMC extends the approaches by Paez et al. (2008), Cai et al. (2013) and Bakar et al. (2015) since it includes a space-time interaction term in the intercept, thus enabling control for complex spatio-temporal confounding patterns.
As a fourth major advantage, following Greven et al. (2011) and Janes et al. (2007), the exposure effect is decomposed into two orthogonal components, each describing the (non-linear) association with the health outcome at a different scale of variation (see Section 3). The focus of inference is the component that captures the association at a finer scale — where confounding by unmeasured factors is less likely (Greven et al., 2011) — which is assumed to exhibit spatio-temporal dynamics. Furthermore, SDGLMC provides explicit interpretation of each part of these dynamics through an analysis-of-variance-like decomposition (see Section 3.1).
This paper is organized as follows. Section 2 offers a brief description of the data that motivated this study. Sections 3 and 4 provide detailed information on the proposed model and on the inferential procedure, respectively. Section 5 presents existing approaches from the literature to which the results of the proposed model are compared. In Section 6, a simulation study is conducted to investigate the ability of SDGLMC in controlling for spatio-temporal confounding when the true association is known. The results from the real-data application are analyzed in Section 7. Finally, Section 8 provides a discussion.
2 The Data
@@
This section is dedicated to the description of the dataset that motivated our research, specifically we analyze the association between daily PM2.5 concentrations and all-cause mortality counts among the elderly (i.e. people aged 65 or more) in Piemonte and Lombardia regions (Italy) between 2018 and 2022.
@@
The data are retrieved from publicly available sources, and full details are given in Section A of the Supplementary Material. Mortality counts are collected from the Italian National Institute of Statistics (ISTAT, 2023a), and are observed on an irregular lattice comprising ${N}=117$ non-overlapping areal units termed health districts (HDs). The study period comprises ${T}=1826$ consecutive days, hence $Y_{it}$ denotes the number of deaths (by any cause) occurred in district $i=1,\dots,{N}$ during day $t=1,\dots,{T}$. The PM2.5 concentrations are used as a measure of the exposure in district $i$ during day $t$, denoted as $X_{it}$, and are collected from the Copernicus Atmosphere Monitoring Service’s (CAMS) Atmosphere Data Store (INERIS et al., 2022, CAMS, 2023). We also acquire information on temperature and relative humidity, which are potential confounders, from the ERA5 reanalyses dataset available at the Copernicus Climate Change Service’s (C3S) Climate Data Store (C3S, 2023, Hersbach et al., 2018). The remote-sensed measurements are spatially and temporally aggregated to match the resolution of the health data.
@@
Figure 2 shows the average daily SMR (on the logarithmic scale) for Lombardia and Piemonte. The seasonal pattern is similar in both regions, with the lowest values in summer, except for the year 2020. Two peaks can be noticed, where the log(SMR) more than doubled: the first one occurred in April 2020 and the other one in November 2020. These are due to the first and second “waves” of the SARS-CoV-2 pandemic, respectively. The first wave was more severe in Lombardia, while the second one was more severe in Piemonte. In addition, the two peaks in the SMR in 2020 are not reflected in the PM2.5 time series. As can be noticed from Table B.3 in the Supplementary Material, the overall 2020 annual mean of PM2.5 levels was the lowest among the five years considered in our analysis.
@@
Figure 2: Time series of the average daily SMR (on the log scale, orange lines) and average daily PM2.5 concentrations (black lines) for Piemonte (top) and Lombardia (bottom) regions. The vertical dashed lines separate each calendar year. Blue vertical bands correspond to the national lockdown periods.
@@
Figure 2 also highlights other two aspects. One is that the highest values of the SMR in 2018 were reached at the beginning of January, suggesting that the influenza epidemic reached its peak before the end of 2017. This is different from the year 2019, where the SMR peaks were reached in February. The other aspect is that there was an increase in mortality during summer 2022. Possible reasons may be the occurrence of prolonged heat waves (indeed, the average daily temperature time-series has its absolute maximum in July 2022), or the outbreak of an influenza epidemic.
However, to study the association of interest, it is necessary to account for the effects of measured and unmeasured confounders. Failing to do so may lead to incorrect inferences, due to concurvity. The proposed approach, which is detailed in the next section, is framed to overcome this issue, among others.
@@ 3 The Proposed Model
It is assumed that the outcome, $Y_{it}$, follows a Poisson distribution with mean $\mu_{it}$, for $i=1,\dots,{N}$ and $t=1,\dots,{T}$. The linear predictor, $\vartheta_{it}=\ln(\mu_{it})$, is modeled as a function of exposure, $X_{it}$, measured confounders, $\mathbf{M}_{it}=({M}_{1it},\dots,{M}_{{p_{m}}it})^{\prime}$, and an unmeasured confounder, $Z_{it}$, which is a smooth function of space and time. Assuming that ${\mathbf{M}}_{it}$ and $Z_{it}$ are spatially and temporally smoother than the exposure, the model can be specified as follows:
$\vartheta_{it}$ $={\color[rgb]{0,0,0}\ln(\mu_{it})}=o_{it}+\beta_{0}+f(X_{it};\bm{\gamma}_{f})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+u_{it}\,,$ (2)
where $o_{it}=\log(E_{it})$ is the offset, $\beta_{0}$ is an intercept, $f(\cdot;\cdot)$ is an unknown smooth function parameterized by the vector $\bm{\gamma}_{f}$, $\bm{\alpha}$ is a ${p_{m}}$-dimensional vector of coefficients, and $u_{it}$ is a spatially and temporally uncorrelated Gaussian random variable, assumed to have zero mean and constant variance, i.e. $u_{it}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(0,\tau_{u}^{2})$.
@@
$f(X_{it};\bm{\gamma}_{f})=f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+\varpi_{it}\,,$ (3)
where $\widehat{X}_{it}$ denotes the large-scale (or overall) spatio-temporal trend, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the first derivative of $f(\cdot;\bm{\gamma}_{f})$ with respect to its first argument, evaluated at $\widehat{X}_{it}$, and $\varpi_{it}$ denotes the Taylor’s remainder. To calculate the spatio-temporal trend, $X_{it}$ can be regressed on a smooth function of space and time (see also Janes et al., 2007):
${\color[rgb]{0,0,0}X_{it}=h({\bf s}_{i},t;df)+\nu_{it}\,,}$ (4)
where ${\bf s}_{i}$ denotes the spatial coordinates of the centroid of areal unit $i$, $df$ are the degrees of freedom of the smooth term, and $\nu_{it}$ is an uncorrelated Gaussian random variable with zero mean and constant variance. The overall trend is then defined as the fitted values obtained from the model (4), i.e. $\widehat{X}_{it}=\widehat{h}({\bf s}_{i},t;df)$.
Equation (3) provides a decomposition of the exposure effect into two components describing large and small (or local) scales of variation. These components are functions of $\widehat{X}_{it}$ and $X_{it}-\widehat{X}_{it}$, respectively. They are orthogonal because $X_{it}-\widehat{X}_{it}$ is the residual term from the regression in Equation (4). Henceforth, the decomposition above should help to disentangle the effect of interest from that of unmeasured confounding factors. The rationale for using detrended exposure, as discussed in the econometric and spatio-temporal literature, is to remove long-term and large-scale spatial variation, preserving only short-term and local fluctuations. This helps reduce the risk of spurious associations driven by underlying trends (Wooldridge, 2020). From an epidemiological point of view, consider that $\widehat{X}_{it}$ captures trend and seasonal components of the exposure, therefore it is expected to be highly correlated with $Z_{it}$, given that the latter is smooth, while the correlation between $X_{it}-\widehat{X}_{it}$ and $Z_{it}$ should be close to zero. Therefore, the unmeasured confounder is more likely to affect the association between outcome and exposure at large scales (Greven et al., 2011, Janes et al., 2007), so the first term in the decomposition of $f(X_{it};\bm{\gamma}_{f})$ in Equation (3) might be biased, but it is less likely that the second term, $f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})$, is affected by confounding bias.
@@
$\vartheta_{it}=o_{it}+\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+Z_{it}+\tilde{u}_{it}\,.$ (5)
@@
For each $i$, the unknown function evaluated at the overall trend of the exposure, $f(\widehat{X}_{it};\bm{\gamma}_{f})$, and its first derivative in Equation (5) can be seen as functions of time, hence time-varying coefficients can adequately represent them (see also West and Harrison, 1997, Section 3.1). Therefore, we propose to replace the linear predictor in Equation (2) with:
$\vartheta_{it}=o_{it}+\beta_{0it}+\beta_{1it}(X_{it}-\widehat{X}_{it})+{\mathbf{M}}_{it}^{\prime}\bm{\alpha}+\tilde{u}_{it}\,,$ (6)
where $\beta_{1it}=f^{\prime}(\widehat{X}_{it};\bm{\gamma}_{f})$ is the spacetime-varying coefficient representing the local association between exposure and outcome (in the sense of Janes et al. (2007)), while $\beta_{0it}$ is a spacetime-varying intercept,
$\beta_{0it}=\beta_{0}+f(\widehat{X}_{it};\bm{\gamma}_{f})+Z_{it}\,,$
@@
We now provide two examples of the function $f(\cdot;\cdot)$ and of the resulting varying coefficients in Equation (6). These show that our model generalizes other methods in the literature.